OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 2x – 3y = 6? it a fraction of some sort.

OpenStudy (anonymous):

If you make this into y=mx+b form, it becomes: 3y=2x-6 --> y=2/3x-2. Parallel lines have the same slope as the original line, so the parallel slope would be 2/3!